Question 29

If a demand curve and a supply curve can be stated functionally as QD = 100 - 5p; and Qs = 90 + 5p, respectively, then the

equilibrium quantity and price would be


1
Expert's answer
2021-10-05T13:53:47-0400

the correct answer is 95 and 1.

At equilibrium, demand = supply

100-5P = 90 + 5P

10 = 10P implies P = 1 and Q = 100 - 5(1) = 95